OpenStudy (anonymous):

help me plzzzzzz @pepperonipizza Solve. x – 7 = –19 A. –26 B. –12 C. 12 D. 26

OpenStudy (anonymous):

@Rina.r can u help me

OpenStudy (anonymous):

and u too @pepperonipizza

OpenStudy (rina.r):

its B x=-12

OpenStudy (anonymous):

ty @pepperonipizza is she right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i have 4 more

OpenStudy (rina.r):

x – 7 = –19 x=-19+7 x=-12

OpenStudy (anonymous):

3 = –(–y + 6) A. –9 B. –3 C. 3 D. 9

OpenStudy (rina.r):

bring it on i will try my best to help.

OpenStudy (anonymous):

omg ty

OpenStudy (rina.r):

its -12 @haseeb96

OpenStudy (haseeb96):

y=3+6 y=9 this is the answer

OpenStudy (anonymous):

omg whitch one is right

OpenStudy (anonymous):

x-7=-19 x=-19+7 x=-12 so therefore....B

OpenStudy (anonymous):

this one i just put up

OpenStudy (anonymous):

3 = –(–y + 6) A. –9 B. –3 C. 3 D. 9

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

im like super confused

OpenStudy (rina.r):

for 2 its B -3

OpenStudy (haseeb96):

x -7 =-19 x=-19+7 x=-12 B is the correct answer

OpenStudy (anonymous):

ty i have 2 more

OpenStudy (anonymous):

Solve for x. |x| – 4 = 9 A. x = 5 or –5 B. x = 5 C. x = 13 or –13 D. x = 13

OpenStudy (haseeb96):

D is the correct answer Africa2

OpenStudy (rina.r):

for question 1 its x=-12 for question2 y=-3

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

on the absolute value question. -13 and 13 are the same value while inside absolute values.

OpenStudy (rina.r):

for question 3 D is teh answer 13

OpenStudy (anonymous):

whitch one is right for 2

OpenStudy (rina.r):

-3 for 2

OpenStudy (haseeb96):

x =13 is the correct answer I mean D

OpenStudy (anonymous):

he said D she said B

OpenStudy (anonymous):

heres question 4

OpenStudy (anonymous):

Serena parachutes from a plane at a height of 20,000 ft above sea level. Several minutes later she notes from an instrument on her wrist that she had dropped 15,000 ft. How far above sea level is she now? A. 5000 B. 10,000 C. –5000 D. –15,000

OpenStudy (rina.r):

guyz are we having a competition like seriously she asked me for help not all of u.

OpenStudy (haseeb96):

D is correct i am sure for it 100 %

OpenStudy (anonymous):

im not a girl im a boy

OpenStudy (haseeb96):

A is correct anwer

OpenStudy (anonymous):

ty for question 4

OpenStudy (rina.r):

yeah A is correct. 5000

OpenStudy (haseeb96):

hahahaha

OpenStudy (anonymous):

heres question 5

OpenStudy (rina.r):

ok

OpenStudy (rina.r):

bring it on

OpenStudy (anonymous):

The variable t represents the temperature at 4 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Mansfield dropped 17°F between 4 p.m. and 11 p.m. The temperature at 11 p.m. was –11°F. What was the temperature at 4 p.m.? A. t – 17 = –11 B. 17 + 11 = t C. 11 – 17 = t D. 17 – t = –11

OpenStudy (rina.r):

its C b coz u -

OpenStudy (haseeb96):

at 4 p.m. A is correct

OpenStudy (anonymous):

hold up C OR B

OpenStudy (rina.r):

THEY ARE ASKING FOR 4 P.M TEMP... UNKNOWN?

OpenStudy (anonymous):

hold up wat grade r u all in

OpenStudy (rina.r):

its C

OpenStudy (anonymous):

im in 8th

OpenStudy (haseeb96):

i am in 12 th

OpenStudy (haseeb96):

its A

OpenStudy (anonymous):

A OR C

OpenStudy (rina.r):

where are u from haseeb?

OpenStudy (anonymous):

im fron gary

OpenStudy (anonymous):

so is it C OR B

OpenStudy (haseeb96):

the temp: at 11 p.m. is -11F the difference between 4 and 11 is 17 F So A is correct t-17=-11 t=-11+17 t=6F

OpenStudy (haseeb96):

At 4 p.m.

OpenStudy (anonymous):

kk

OpenStudy (haseeb96):

i am from Hyderabad and u?

OpenStudy (anonymous):

yes i passed

OpenStudy (anonymous):

were is Hyderabad

OpenStudy (anonymous):

@Haseeb96 can u help me some more plzzz

OpenStudy (haseeb96):

yes. if i did then i will help u Say rina she wil also help u

OpenStudy (anonymous):

kkk @Rina.r r u still going to help me

OpenStudy (anonymous):

werre is Hyderabad at

OpenStudy (haseeb96):

in Pakistan

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Evaluate x + y when x = 14.4 and y = 3.1 A. 45.5 B. 17.5 C. 11.3 D. –17.5

OpenStudy (haseeb96):

B is correct answer

OpenStudy (anonymous):

kk how u get the answer

OpenStudy (haseeb96):

x+y=14.4+3.1 since , put the given values of x and y in the given Equation =17.5

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

hers the other 11

OpenStudy (anonymous):

Evaluate a + b when a = –16.2 and b = –11.4 A. –27.6 B. –4.8 C. 4.8 D. 27.6

OpenStudy (haseeb96):

please call the Rina

OpenStudy (anonymous):

i did shes here but not answering

OpenStudy (haseeb96):

A is Correct

OpenStudy (anonymous):

ty hers the other1

OpenStudy (rina.r):

hey i am sorry i was away

OpenStudy (anonymous):

Write an expression to model the following situation: Maria and Seth are baking cookies for the annual bake sale. Maria has baked 58 cookies, and Seth has baked 117. How many cookies have they baked in all? A. 58 + 117 B. 58 – 117 C. 117 – 58 D. 117 + (–58)

OpenStudy (haseeb96):

Boy , just make your mind sharp and solve the problems without any hesitations

OpenStudy (haseeb96):

A is the correct answer

OpenStudy (anonymous):

ty heres the other 1

OpenStudy (anonymous):

One day, the temperature started at 14 degrees at 6:00 a.m., then climbed 5 degrees by noon, and then dropped 7 degrees by midnight. What was the temperature at midnight? A. 2 degrees B. 12 degrees C. 16 degrees D. 26 degrees

OpenStudy (haseeb96):

B is the correct answer

OpenStudy (rina.r):

for cookies question its A

OpenStudy (anonymous):

omg here we go again

OpenStudy (anonymous):

A or B

OpenStudy (rina.r):

its B

OpenStudy (anonymous):

heres the other 1

OpenStudy (rina.r):

OK

OpenStudy (anonymous):

What is the missing reason in the proof? Prove –(–y – x) – x = y Statements Reasons –(–y – x) – x = –[–y +(–x)] – x Definition of subtraction –[–y +(–x)] – x = y + x – x Opposite of a sum property y + x – x = y + x + (–x) Definition of subtraction y + x + (–x) = y + [x + (–x)] Associative Property of Addition y + [x + (–x)] = y + 0 Additive Inverse Property y + 0 = y ________________________ A. Symmetric Property B. Additive Inverse Property C. Additive Identity Property D. Opposite of a Sum Property

OpenStudy (rina.r):

I THINK A

OpenStudy (anonymous):

so u think A u not positive